OpenStudy (anonymous):

Nina is 10 years younger than Deepak. Deepak is 3 times as old as Nina. Which system of equations can be used to find d, Deepak's age, and n, Nina's age?

OpenStudy (anonymous):

answer choices?

OpenStudy (anonymous):

d = n + 10 d = 3n d = n − 10 d = 3n d = n + 10 n = 3d d = n − 10 n = 3d

OpenStudy (anonymous):

the boxes are -

OpenStudy (anonymous):

i dont know why it dose that when i copy and paste

OpenStudy (anonymous):

its ok do you have an idea of which one it is?

OpenStudy (anonymous):

i think it is c

OpenStudy (anonymous):

im confusin myself lol so this might take a bit of time

OpenStudy (anonymous):

i think its b because it says she is ten years younger so youll have to minus her age by ten to get that then it says he is three times that age so youll have too do 3 times ninas age but i could defiantly be wrong so go with what seems correct